OpenStudy (majikdusty):

Find the local extrema of f (x^2)/sqrt(x+7)

OpenStudy (helder_edwin):

u have \[ \large f(x)=\frac{x^2}{\sqrt{x+7}} \] find f'(x) (the first derivative)

OpenStudy (majikdusty):

I got \[(2x \sqrt{x+7}-(x^2)/2)\div((x+7)\sqrt{x+7})\]

OpenStudy (majikdusty):

is this correct?

OpenStudy (helder_edwin):

\[ \large f'(x)=\frac{2x\sqrt{x+7}-x^2\frac{1}{2\sqrt{x+7}}}{x+7} \]